# ElastiBench: Scalable Continuous Benchmarking on Cloud FaaS Platforms

### v1

@@ 1. Introduction
Microbenchmarks offer a fine-grained insight into the performance of an application by repeatedly benchmarking a small part of the application, e.g., on the function or package level (Laaber et al., 2021a). This allows developers to gain detailed insight into the impact their code changes have on specific code paths. To achieve this, microbenchmarks can be run as part of a CI/CD (continuous integration and deployment) pipeline, which lets developers detect and fix performance regressions before they are introduced to the live environment (Everett Pompeii, 2024; Anzt et al., 2019; Pearce et al., 2023).
@@
Current approaches of executing microbenchmarks in cloud environments rely on virtual machines, which requires multiple repetitions of every microbenchmark to achieve reliable and statistically significant results (Grambow et al., 2023a; Laaber et al., 2019; Bermbach et al., 2017). For example, in previous work on microbenchmarking in open-source code bases, we needed to repeat benchmarks more than 40 times on different virtual machines to achieve reliable results. This took ~4h and cost ~\$1.14 for VictoriaMetrics, [^1] https://github.com/VictoriaMetrics/VictoriaMetrics and took ~11h and cost ~\$3.15 for InfluxDB [^2] https://github.com/influxdata/influxdb (Grambow et al., 2023a). These benchmark durations and costs are prohibitive for integrating performance regression detection in CI/CD pipelines for every release (let alone every commit), where continuous evaluation with rapid developer feedback is paramount.
@@
Figure 1. The traditional approach of microbenchmarking (top) relies on executing tasks (different circles) in random order multiple times on different virtual machines (gray boxes) to get reliable results. Using FaaS, these tasks can be executed on multiple function instances in parallel (bottom). With instance parallelism (three in this example), the duration of the suite run can be drastically reduced while also reducing inter-microbenchmark influences.
@@
We propose leveraging Function-as-a-Service (FaaS) to massively scale out microbenchmark execution with little lead time, as shown in Figure 1: The elastic scalability of cloud FaaS platforms allows us to run hundreds of microbenchmarks in parallel, achieving robust results in a fraction of the time sequential microbenchmark execution requires. This enables developers to run them often as a normal part of their development workflows, e.g., to evaluate whether a local change has performance implications.
@@
The key challenge in deploying microbenchmarks on FaaS is the unique execution environment that is prone to performance variability through cold-starts, temporal variability, and infrastructure heterogeneity (Bermbach et al., 2020; Schirmer et al., 2023). With ElastiBench, we overcome this by focusing on the detection of relative performance regression between two versions of the same software: For this, we deploy both versions in the same FaaS function instance (similar to duet benchmarking (Bulej et al., 2020)), thus ensuring an identical execution environment. We then repeatedly call this function, triggering multiple microbenchmarks runs (see also Figure 1). By carefully managing the execution order of microbenchmarks and their targets as well as the statistical analysis process afterwards, we can mitigate the noise effects of the FaaS platform and maintain the robust detection capabilities of the microbenchmark suite. At the same time, we gain the significant speed-up and cost reductions of the FaaS environment, making our approach far superior to the state-of-the-art. In this regard, we present the following contributions:
  • We analyze the main challenges in microbenchmarking using FaaS (Section 3).
@@
  • We propose ElastiBench, an approach for executing microbenchmark suites on cloud-based FaaS platforms (Section 4).
@@
  • We implement a proof-of-concept prototype of ElastiBench for Go microbenchmarks on AWS Lambda (Section 5).
@@
  • We evaluate the detection capabilities and the efficiency of ElastiBench by comparing it to published results that used cloud VMs (Section 6).
2. Microbenchmarking in the Cloud
@@
Microbenchmarks are short-running performance tests that evaluate a Software Under Test (SUT) on a smaller scale than full application benchmarks (Grambow et al., 2023a). Their scope can be compared to that of unit tests, i.e., they measure the performance of a single method or package (Laaber et al., 2019). [^3] To disambiguate from FaaS functions, we refer to the component that is being evaluated as methods, independently of what they are called in their respective programming language. Previous research has shown that microbenchmarks are easier to set up than application benchmarks, which can require other components to run (Grambow et al., 2021a). In compiled languages, the corresponding microbenchmark test files are compiled and then repeatedly executed. In the Go programming language, for example, every method that is in a file with the suffix _test.go and starts with Benchmark counts as a microbenchmark and can be executed using the go test command line tool (The Go Authors, 2024). With a configurable overall duration, the tool then reports the total amount of iterations that were executed, and the average time per execution. Usually, executing a single benchmark takes on the order of a few seconds (Laaber et al., 2019; Cheney, 2019).
@@
Especially in cloud environments using shared resources, the total duration of a microbenchmark varies between different runs. This is due to the inherent variability of the code that is executed, and effects within as well as between instances (Grambow et al., 2023a). When trying to measure the application performance, the goal is to minimize the impact of instance variability to capture the true performance of the benchmark. Thus, the goal is not to achieve a consistent result, as the microbenchmark itself might be highly variable, especially if it is written in an interpreted language (Laaber et al., 2019). To limit measurement error, the order in which microbenchmarks are executed is randomized (Randomized Multiple Interleaved Trials, RMIT (Abedi and Brecht, 2017)) so that order-effects are averaged out. Next, to limit the impact of performance variance between instances, the two different versions of the SUT are executed on the same virtual machine and only their relative difference is taken into account for performance measurements. These experiments are then repeated on different virtual machines (Grambow et al., 2023a). Current research indicates that 5 to 30 iterations of a microbenchmark are enough to get reliable results (Laaber et al., 2019).
@@
Developers use microbenchmarks to detect whether a code change had performance implications. Thus, microbenchmarking results from the previous and new version of the software are collected and analyzed to detect statistically significant changes. Current research uses the median performance change to measure the difference between two versions of a microbenchmark because it is more robust to outliers and skewed data than averages (Maricq et al., 2018). A standard methodology to calculate the confidence interval of the median is bootstrapping, which we also used in our previous work (Kalibera and Jones, 2020; Grambow et al., 2023a; Bulej et al., 2017b, a; SciPy Community, 2024; Maricq et al., 2018). Bootstrapping randomly resamples from the population with replacement to generate multiple bootstrap samples and calculates their median. These random samples are then used to calculate confidence intervals around the measured median. According to Laaber et al. (Laaber et al., 2019), the underlying performance variability in microbenchmarks executed in the cloud is so high that a small performance change (3% to 10% in different studies(Georges et al., 2007; Mytkowicz et al., 2009)) does not necessarily imply a real-world change in performance even if it is statistically significant.
3. Challenges for Microbenchmarking on FaaS
@@
FaaS platforms enable developers to easily create massively scalable applications. The main use case for serverless platforms are short-running, resource-light functions that take in the order of seconds to complete (Eismann et al., 2021). FaaS platforms are an efficient execution environment for such tasks as they can benefit from bin-packing different small function executions from different tenants on shared hardware. Cloud FaaS platforms in particular rely on optimized execution environments to quickly spin up function instances on demand, i.e., cold starts, and by overprovisioning resources (Manner et al., 2018; Agache et al., 2020; Brooker et al., 2023). This has interesting implications for using FaaS platforms to run microbenchmarks, which are sensitive to performance variability given their short execution duration.
3.1. Performance
@@
Due to their small resource footprint, the FaaS provider can run many function instances on the same computing resources. The smallest functions on Google Cloud Functions and AWS Lambda, for example, have access to less than 10% of one vCPU (Cordingly et al., 2022), which leads to high interference between functions sharing compute resources. Previous research has shown that function performance can vary up to 15% diurnally and, even if executed at the same time, the variation between function instances can be considerable (Schirmer et al., 2023). These performance variations increase the uncertainty of FaaS-based microbenchmarking results, as it is unknown to which degree an experiment has been affected by underlying performance changes. Additionally, the performance difference between a cold and warm start can also influence benchmarking results, as during a cold start the application code might not be optimized yet, and some parts of the file system might be lazily loaded (Brooker et al., 2023).
3.2. Restricted Environment
@@
3.3. Representative Environment
When executing microbenchmarks, the experiment environment should resemble the real use case as much as possible to achieve relevant results (Bermbach et al., 2017). FaaS platforms offer a standardized interface with only a few configuration options, which might be limiting for use cases with unique resource demands. Additionally, while the platforms allow developers to choose a processor architecture, the exact CPU model used is not configurable and varies even in one availability zone of a provider (Cordingly et al., 2023).
4. Executing Microbenchmarks in FaaS
Figure 2. The process used to collect benchmarking data from FaaS functions. First, the function image (cf. Section 5) is built and deployed to the FaaS platform. In step two, the function is called repeatedly with configurable repeats per microbenchmarks and instance parallelism. The results of all calls are then analyzed on the calling machine. The calling system can be the workstation of a developer or an automated CI/CD pipeline. While this figure shows multiple microbenchmarks being executed in one function call, Figure 1 shows the extreme example of just one microbenchmark per function call.
@@ 4. Executing Microbenchmarks in FaaS
Execution of microbenchmarks is an embarrassingly parallelizable workload, since they have no dependencies on each other. This makes them well suited as an application of FaaS. Furthermore, two challenges of traditional microbenchmarking that are not as relevant on FaaS are execution-order effects, which might falsify results when microbenchmarks are executed in sequence on a single node, and the performance variability between a low number of virtual machine instances (Grambow et al., 2023a; Laaber et al., 2019). When using FaaS functions with high parallelism, the impact of a function instance with outlier performance is lower due to the sheer number of instances. Additionally, the execution order of benchmarks within a function instance can be randomized as well: by only executing one microbenchmark per function invocation, the platform assigns every microbenchmark to an instance using its own control logic, which means that the order effects are averaged out with sufficiently many repeats.
@@
Otherwise, the same limitations that apply when running microbenchmarks on cloud-based virtual machines (Grambow et al., 2023a; Laaber et al., 2019) apply to FaaS as well: Due to the inherent performance variability of cloud platforms, a microbenchmark should always be run for each version of the software within the same function instance. In the statistical analysis, only the relative performance change between these versions is taken into account. This relative change is transitive, i.e., subsequent versions of the software only need to be compared to a reference version and their relative difference to other versions can be calculated based on that.
@@
Figure 2 shows the process for running microbenchmarks on a FaaS. The runner that is used to conduct performance benchmarks (e.g., the workstation of a developer or a CI/CD runner) first builds the function image that will be called. The image contains the two software versions and the code required to run microbenchmarks (cf. Section 5). Afterwards, the newly deployed function is called by the runner with the following parameters: the microbenchmarks that should be executed, whether their order and the order in which to run the different SUT versions should be randomized, and timeouts for the benchmark runs. The microbenchmark execution can be repeated many times in different function calls. This way, the system takes advantage of the transparent assignment of calls to function instances: by randomizing the order of function calls, the instances on which the repeats of microbenchmarks are executed is also effectively randomized. After all microbenchmarks have been executed sufficiently often, the function is obsolete and can be deleted: since the SUT source is part of the function image and the comparison of these two versions has been finished, the function will not be called again. All other container layers of the function image (e.g., the actual function) can be reused for future experiments. Depending on the number of microbenchmarks and the instance parallelism, the whole process can complete within the maximum time the FaaS platform keeps a function instance alive (e.g., ~15min for AWS Lambda and Google Cloud Functions). Since the short overall duration of a benchmark run essentially means that only the first requests will be cold starts, the runner can influence the total duration and total cost of the benchmarking run by configuring the amount of requests that are sent in parallel: higher parallelism leads to shorter runs, while increasing cost due to the increased number of cold starts.
5. Proof-of-Concept Implementation
@@
In this section, we present the prototype function we implemented to execute microbenchmarks with the ElastiBench approach. We focus on SUTs written in Go, as it is a popular programming language for cloud systems and has native support for running microbenchmarks. We implement our prototype for AWS Lambda since it is one of the most mature and popular platforms and showed comparably low performance variation in previous research (Schirmer et al., 2023; Copik et al., 2021). The architecture and prototype do not require any features that are specific to AWS Lambda, so that it could easily be adapted for other platforms. We describe the components comprising the function image in the following sections. The total file size of an image for a complex SUT can be bigger than ~1GB, of which ~240MB are components that are necessary for the microbenchmarks to work, and the rest is taken up by the SUT as well as a prepopulated Go build cache of the SUT source code.
@@
To speed up repeated compilation in benchmarks, the microbenchmarking pipeline uses a build cache that stores precompiled binaries. Without a prepopulated cache, cold starts would take even longer as the cache would have to be filled during the first call. Thus, we fill this cache during the build process (i.e., on the machine of the developer). While this increase in image size (almost 1GB for the project used in our evaluation) also increases cold start times since more data needs to be transferred to function runners, it still leads to overall faster and more consistent performance. Due to caching in function runners (Brooker et al., 2023), the first cold starts after deploying a new version of the function are slower, but subsequent cold starts benefit from caching of these larger images.
@@
Due to the way that FaaS providers limit writable locations in the file system (see also Section 3), the prepopulated cache location is not writable during function execution. To work around this, Go supports custom cacher implementations (bradfitz, 2024). In the function image, a custom cacher is used to read the prepopulated cache but writes changes to another (writable) directory. When the microbenchmarking pipeline tries to access cached binaries, the custom cacher first checks the instance cache and then falls back to the prepopulated cache. The file size of the custom cacher program is ~3MB.
6. Evaluation
In this section, we present how the experiments are designed (Section 6.1), their results (Section 6.2), and discuss key takeaways (Section 6.3).
6.1. Experiment Design
The following sections contain descriptions of the SUT used in our evaluation (Section 6), the specific methods we use for our analysis (Section 6), and give an overview of the experiments (Section 6).

### v2

@@ I Introduction
Microbenchmarks offer a fine-grained insight into the performance of an application by repeatedly benchmarking a small part of the application, e.g., on the function or package level [1]. This allows developers to gain detailed insight into the impact their code changes have on specific code paths. To achieve this, microbenchmarks can be run as part of a CI/CD (continuous integration and deployment) pipeline, which lets developers detect and fix performance regressions before they are introduced to the live environment [2, 3, 4].
@@
Current approaches of executing microbenchmarks in cloud environments rely on virtual machines, which requires multiple repetitions of every microbenchmark to achieve reliable and statistically significant results [5, 6, 7]. For example, in previous work on microbenchmarking in open-source code bases, we needed to repeat benchmarks more than 40 times on different virtual machines to achieve reliable results. This took ~4h and cost ~\$1.14 for VictoriaMetrics, [^1] https://github.com/VictoriaMetrics/VictoriaMetrics and took ~11h and cost ~\$3.15 for InfluxDB [^2] https://github.com/influxdata/influxdb [5]. These benchmark durations and costs are prohibitive for integrating performance regression detection in CI/CD pipelines for every release (let alone every commit), where continuous evaluation with rapid developer feedback is paramount.
@@
Figure 1: The traditional approach of microbenchmarking (top) relies on executing tasks (different circles) in random order multiple times on different virtual machines (gray boxes) to get reliable results. Using FaaS, these tasks can be executed on multiple function instances in parallel (bottom). With instance parallelism (three in this example), the duration of the suite run can be drastically reduced while also reducing inter-microbenchmark influences.
@@
We propose leveraging Function-as-a-Service (FaaS) to massively scale out microbenchmark execution with little lead time, as shown in Fig. 1: The elastic scalability of cloud FaaS platforms allows us to run hundreds of microbenchmarks in parallel, achieving robust results in a fraction of the time sequential microbenchmark execution requires. This enables developers to run them often as a normal part of their development workflows, e.g., to evaluate whether a local change has performance implications.
@@
The key challenge in deploying microbenchmarks on FaaS is the unique execution environment that is prone to performance variability through cold-starts, temporal variability, and infrastructure heterogeneity [8, 9]. With ElastiBench, we overcome this by focusing on the detection of relative performance regression between two versions of the same software: For this, we deploy both versions in the same FaaS function instance (similar to duet benchmarking [10]), thus ensuring an identical execution environment. We then repeatedly call this function, triggering multiple microbenchmarks runs (see also Fig. 1). By carefully managing the execution order of microbenchmarks and their targets as well as the statistical analysis process afterwards, we can mitigate the noise effects of the FaaS platform and maintain the robust detection capabilities of the microbenchmark suite. At the same time, we gain the significant speed-up and cost reductions of the FaaS environment, making our approach far superior to the state-of-the-art. In this regard, we present the following contributions:
  • We analyze the main challenges in microbenchmarking using FaaS (Section III).
@@
  • We propose ElastiBench, an approach for executing microbenchmark suites on cloud-based FaaS platforms (Section IV).
@@
  • We implement a proof-of-concept prototype of ElastiBench for Go microbenchmarks on AWS Lambda (Section V).
@@
  • We evaluate the detection capabilities and the efficiency of ElastiBench by comparing it to published results that used cloud VMs (Section VI).
II Microbenchmarking in the Cloud
@@
Microbenchmarks are short-running performance tests that evaluate a Software Under Test (SUT) on a smaller scale than full application benchmarks [5]. Their scope can be compared to that of unit tests, i.e., they measure the performance of a single method or package [6]. [^3] To disambiguate from FaaS functions, we refer to the component that is being evaluated as methods, independently of what they are called in their respective programming language. Previous research has shown that microbenchmarks are easier to set up than application benchmarks, which can require other components to run [11]. In compiled languages, the corresponding microbenchmark test files are compiled and then repeatedly executed. In the Go programming language, for example, every method that is in a file with the suffix _test.go and starts with Benchmark counts as a microbenchmark and can be executed using the go test command line tool [12]. With a configurable overall duration, the tool then reports the total amount of iterations that were executed, and the average time per execution. Usually, executing a single benchmark takes on the order of a few seconds [6, 13].
@@
Especially in cloud environments using shared resources, the total duration of a microbenchmark varies between different runs. This is due to the inherent variability of the code that is executed, and effects within as well as between instances [5]. When trying to measure the application performance, the goal is to minimize the impact of instance variability to capture the true performance of the benchmark. Thus, the goal is not to achieve a consistent result, as the microbenchmark itself might be highly variable, especially if it is written in an interpreted language [6]. To limit measurement error, the order in which microbenchmarks are executed is randomized (Randomized Multiple Interleaved Trials, RMIT [14]) so that order-effects are averaged out. Next, to limit the impact of performance variance between instances, the two different versions of the SUT are executed on the same virtual machine and only their relative difference is taken into account for performance measurements. These experiments are then repeated on different virtual machines [5]. Current research indicates that 5 to 30 iterations of a microbenchmark are enough to get reliable results [6].
@@
Developers use microbenchmarks to detect whether a code change had performance implications. Thus, microbenchmarking results from the previous and new version of the software are collected and analyzed to detect statistically significant changes. Current research uses the median performance change to measure the difference between two versions of a microbenchmark because it is more robust to outliers and skewed data than averages [15]. A standard methodology to calculate the confidence interval of the median is bootstrapping, which we also used in our previous work [16, 5, 17, 18, 19, 15]. Bootstrapping randomly resamples from the population with replacement to generate multiple bootstrap samples and calculates their median. These random samples are then used to calculate confidence intervals around the measured median. According to Laaber et al. [6], the underlying performance variability in microbenchmarks executed in the cloud is so high that a small performance change (3% to 10% in different studies[20, 21]) does not necessarily imply a real-world change in performance even if it is statistically significant.
III Challenges for Microbenchmarking on FaaS
@@
FaaS platforms enable developers to easily create massively scalable applications. The main use case for serverless platforms are short-running, resource-light functions that take on the order of seconds to complete [22]. FaaS platforms are an efficient execution environment for such tasks as they can benefit from bin-packing different small function executions from different tenants on shared hardware. Cloud FaaS platforms in particular rely on optimized execution environments to quickly spin up function instances on demand, i.e., cold starts, and by overprovisioning resources [23, 24, 25]. This has interesting implications for using FaaS platforms to run microbenchmarks, which are sensitive to performance variability given their short execution duration.
III-A Performance
@@
Due to their small resource footprint, the FaaS provider can run many function instances on the same computing resources. The smallest functions on Google Cloud Functions and AWS Lambda, for example, have access to less than 10% of one vCPU [26], which leads to high interference between functions sharing compute resources. Previous research has shown that function performance can vary up to 15% diurnally and, even if executed at the same time, the variation between function instances can be considerable [9]. These performance variations increase the uncertainty of FaaS-based microbenchmarking results, as it is unknown to which degree an experiment has been affected by underlying performance changes. Additionally, the performance difference between a cold and warm start can also influence benchmarking results, as during a cold start the application code might not be optimized yet, and some parts of the file system might be lazily loaded [25].
III-B Restricted Environment
@@
III-C Representative Environment
When executing microbenchmarks, the experiment environment should resemble the real use case as much as possible to achieve relevant results [7]. FaaS platforms offer a standardized interface with only a few configuration options, which might be limiting for use cases with unique resource demands. Additionally, while the platforms allow developers to choose a processor architecture, the exact CPU model used is not configurable and has been shown to vary within individual availability zones of a provider [27].
IV Executing Microbenchmarks in FaaS
Figure 2: The process used to collect benchmarking data from FaaS functions. First, the function image (cf. Section V) is built and deployed to the FaaS platform. In step two, the function is called repeatedly with configurable repeats per microbenchmarks and instance parallelism. The results of all calls are then analyzed on the calling machine. The calling system can be the workstation of a developer or an automated CI/CD pipeline. While this figure shows multiple microbenchmarks being executed in one function call, Fig. 1 shows the extreme example of just one microbenchmark per function call.
@@ IV Executing Microbenchmarks in FaaS
Execution of microbenchmarks is an embarrassingly parallelizable workload, since they have no dependencies on each other. This makes them well suited as an application of FaaS. Furthermore, two challenges of traditional microbenchmarking that are not as relevant on FaaS are execution-order effects, which might falsify results when microbenchmarks are executed in sequence on a single node, and the performance variability between a low number of virtual machine instances [5, 6]. When using FaaS functions with high parallelism, the impact of a function instance with outlier performance is lower due to the sheer number of instances. Additionally, the execution order of benchmarks within a function instance can be randomized as well: by only executing one microbenchmark per function invocation, the platform assigns every microbenchmark to an instance using its own control logic, which means that the order effects are averaged out with sufficiently many repeats.
@@
Otherwise, the same limitations that apply when running microbenchmarks on cloud-based virtual machines [5, 6] apply to FaaS as well: Due to the inherent performance variability of cloud platforms, a microbenchmark should always be run for each version of the software within the same function instance. In the statistical analysis, only the relative performance change between these versions is taken into account. This relative change is transitive, i.e., subsequent versions of the software only need to be compared to a reference version and their relative difference to other versions can be calculated based on that.
@@
Fig. 2 shows the process for running microbenchmarks on a FaaS. The runner that is used to conduct performance benchmarks (e.g., the workstation of a developer or a CI/CD runner) first builds the function image that will be called. The image contains the two software versions and the code required to run microbenchmarks (cf. Section V). Afterwards, the newly deployed function is called by the runner with the following parameters: the microbenchmarks that should be executed, whether their order and the order in which to run the different SUT versions should be randomized, and timeouts for the benchmark runs. The microbenchmark execution can be repeated many times in different function calls. This way, the system takes advantage of the transparent assignment of calls to function instances: by randomizing the order of function calls, the instances on which the repeats of microbenchmarks are executed is also effectively randomized. After all microbenchmarks have been executed sufficiently often, the function is obsolete and can be deleted: since the SUT source is part of the function image and the comparison of these two versions has been finished, the function will not be called again. All other container layers of the function image (e.g., the actual function) can be reused for future experiments. Depending on the number of microbenchmarks and the instance parallelism, the whole process can complete within the maximum time the FaaS platform keeps a function instance alive (e.g., ~15min for AWS Lambda and Google Cloud Functions). Since the short overall duration of a benchmark run essentially means that only the first requests will be cold starts, the runner can influence the total duration and total cost of the benchmarking run by configuring the amount of requests that are sent in parallel: higher parallelism leads to shorter runs, while increasing cost due to the increased number of cold starts.
V Proof-of-Concept Implementation
@@
In this section, we present the prototype function we implemented to execute microbenchmarks with the ElastiBench approach. We focus on SUTs written in Go, as it is a popular programming language for cloud systems and has native support for running microbenchmarks. We implement our prototype for AWS Lambda since it is one of the most mature and popular platforms and showed comparably low performance variation in previous research [9, 28]. The prototype is available as open-source [^4] https://github.com/umbrellerde/microbenchmarks-faast. The architecture and prototype do not require any features that are specific to AWS Lambda, so that it could easily be adapted for other platforms. We describe the components comprising the function image in the following sections. The total file size of an image for a complex SUT can be bigger than ~1GB, of which ~240MB are components that are necessary for the microbenchmarks to work, and the rest is taken up by the SUT as well as a prepopulated Go build cache of the SUT source code.
@@
To speed up repeated compilation in benchmarks, the microbenchmarking pipeline uses a build cache that stores precompiled binaries. Without a prepopulated cache, cold starts would take even longer as the cache would have to be filled during the first call. Thus, we fill this cache during the build process (i.e., on the machine of the developer). While this increase in image size (almost 1GB for the project used in our evaluation) also increases cold start times since more data needs to be transferred to function runners, it still leads to overall faster and more consistent performance. Due to caching in function runners [25], the first cold starts after deploying a new version of the function are slower, but subsequent cold starts benefit from caching of these larger images.
@@
Due to the way that FaaS providers limit writable locations in the file system (see also Section III), the prepopulated cache location is not writable during function execution. To work around this, Go supports custom cacher implementations [29]. In the function image, a custom cacher is used to read the prepopulated cache but writes changes to another (writable) directory. When the microbenchmarking pipeline tries to access cached binaries, the custom cacher first checks the instance cache and then falls back to the prepopulated cache. The file size of the custom cacher program is ~3MB.
VI Evaluation
In this section, we present how the experiments are designed (Section VI-A), their results (Section VI-B), and discuss key takeaways (Section VI-C).
VI-A Experiment Design
The following sections contain descriptions of the SUT used in our evaluation (Section VI), the specific methods we use for our analysis (Section VI), and give an overview of the experiments (Section VI).